Three point charges are arranged in horizontal line as shown below. Find the electric force on Q1 given the following: Q1-4 mC, Q2-3 md, Q3-3 mc, r1-80 m, and r2 = 200 m. Remember thatょpositive force points to the right and a negative force points to the left. Q2 Q3 What is the net force on charge Q1? 18.3 (in units of N) Computers answer now shown above. I Tries 2/2 Previous Tries What is the net force on charge Q2? in units of N) Submit Answer Tries 012 What is the net force on charge 03? (in units of N) Submit Answer Tries 0/2 What is the sum of the forces on all three charges? in units of N) Submit Answer Tries 0/2
Answer #1

a] Using Colombs law, Net force on q1 = F1 = - kq1q2/r1^2 - kq1q3/[r1+r3]^2

= -9e9*4e-3*3e-3/80^2 - 9e9*4e-3*3e-3/280^2

= - 18.25 N

b]Using Colombs law, Net force on q2 = F2 = kq1q2/r1^2- kq1q3/[r2]^2

= 9e9*4e-3*3e-3/80^2 - 9e9*3e-3*3e-3/200^2

= 14.85 N

c] Using Colombs law, Net force on q3 = F3 = kq3q2/r2^2+ kq1q3/[r1+r2]^2

= 9e9*3e-3*3e-3/200^2 + 9e9*3e-3*4e-3/280^2

= 3.40 N

d] Fsum = F1+F2+F3 = - 18.25+14.85+3.40 = 0 N